21.) Tris, (HOCH2)3CNH2 (pKb = 5.91), a very common buffer for studying biochemical processes, is prepared by a biological chemist using 10.0M NaOH to a pH of 7.79 using 31.52 g of (HOCH2)3CNH3Cl, or TrisH+, (MW=157.597g/mol, pKa 8.075).

>>>> Need help with second part of question pls>>>>...Now you dilute the solution from problem 21 to 1.0 L and take half of the buffer (500.0 mL) and add 0.0100 moles solid HCl. What is the new pH?
Answer #1

It is a basic buffer bot = pk + log [salt] bot = pkt Base] 14-7.79 = 5.91 + log 310522 157.597 Base = 6.21 = 5.91 + log 0.2 B